Reports 1-1 of 1 Clear search Modify search
Injection system (Laser power stabilization)
derossi, chiummo, cleva - 18:15 Tuesday 05 November 2019 (47547) Print this report
PSTAB glitches

In the last days many glitches seem to come from the PSTAB. We then performed two actions:

- PMC realignment, as already observed and suggested by F. Cleva (#47533).

- PSTAB gain adjustment. We injected a line on the "perturb" input of the PSTAB locking electronics and monitored at the scope the amplitude of the error signals before and after the perturbation. We then adjusted the frequency of the injected line in order to have them of the same amplitude: this frequency is a good estimation of the UGF. It was 120 kHz, and by lowering a bit the gain we brought it at ~ 80 kHz (the corresponding position of the potentiometer is 9 o'clock).  The first plot is a comparison of the current PSTAB err signal (blue) with the one we used to have before Oct 22nd (#47250), when we lowered the power on the photodiodes.

From plot 2 we can observe that the error signal is now much less glitchy, and also from the omicron scan the glitches at 1 kHz seem to have disappear (plot 3).

 

Images attached to this report
Comments to this report:
direnzo - 11:17 Wednesday 06 November 2019 (47557) Print this report

The BNS range drops of the last days seem to be indeed related with the PSTAB, as visible from their correlation obtained with NonNA. In the attached figure, the (anti-)correlation of Hrec_Range_BNS and PSTAB_PD1_AC_FS_rms immediately before the maintenace ofTuesday. The intervention seems to have fixed the problem, and less high frequency glitches and BNS range drops are present now.

Images attached to this comment
Search Help
×

Warning

×